OpenStudy (wow.):

WILL MEDAL!!! Which of the following numbers can be expressed as a decimal that terminates? (A) 7/4 (B) 5/9 (C) 3/8 (D) 2/3

OpenStudy (danjs):

even denominators are nice

OpenStudy (danjs):

try and make each one into a fraction out of 10, or 100 , or 1000, exe...

OpenStudy (danjs):

7/4 = 175/100 = 1 + 75/100 = 1.75

OpenStudy (md152727):

Not B or D for sure

OpenStudy (md152727):

A would be a good choice

OpenStudy (dayakar):

any fraction which has it's denominator in the form of 2^m * 5^m , [ m,n are whole numbers] are terminating decimals remaining are non terminating recurring decimals

OpenStudy (dayakar):

A) \[\frac{ 7 }{ 4 }=\frac{ 7 }{2 ^{2}*5^{0} }\]

OpenStudy (dayakar):

A) is terminating decimal

OpenStudy (dayakar):

do u understand